### v2

# Cartesian atomic cluster expansion for machine learning interatomic potentials

Abstract
Machine learning interatomic potentials are revolutionizing large-scale, accurate atomistic modelling in material science and chemistry. Many potentials use atomic cluster expansion or equivariant message passing frameworks. Such frameworks typically use spherical harmonics as angular basis functions, and then use Clebsch-Gordan contraction to maintain rotational symmetry, which may introduce redundancies in representations and computational overhead. We propose an alternative: a Cartesian-coordinates-based atomic density expansion. This approach provides a complete set of polynormially indepedent features of atomic environments while maintaining interaction body orders. Additionally, we integrate low-dimensional embeddings of various chemical elements and inter-atomic message passing. The resulting potential, named Cartesian Atomic Cluster Expansion (CACE), exhibits good accuracy, stability, and generalizability. We validate its performance in diverse systems, including bulk water, small molecules, and 25-element high-entropy alloys.
I Introduction
Machine learning interatomic potentials (MLIPs) can learn from quantum-mechanical calculations and predict the energy and forces of atomic configurations speedily, thus enabling more precise and comprehensive exploration of material and molecular properties at scale Keith et al. (2021); Unke et al. (2021). Crucially, while atomic Cartesian coordinates encode all the essential information of a structure, they cannot be directly used in learning tasks, due to the lack of symmetries like translation, rotation, inversion, and atom permutations. Consequently, numerous methods have been developed to enforce these symmetries Musil et al. (2021); Unke et al. (2021).
Notably, atomic cluster expansion (ACE) Drautz (2019) represents atomic environments based on a body-ordered expansion, utilizing a complete basis set of spherical harmonics and radial components. ACE can be viewed as a general framework of many other atomic environment representations, such as Atom Centered Symmetry Functions (ACSF) Behler and Parrinello (2007), Smooth Overlap of Atomic Positions (SOAP) descriptor Bartók et al. (2010), Moment Tensor Potentials (MTPs) Shapeev (2016), and bispectrums Bartók et al. (2013).
Another important approach for learning atomic interactions uses message passing neural networks (MPNNs). They represent structures as graphs with atoms as nodes, and apply message passing operations to learn atomic environment representations. Earlier MPNN models, such as SchNet Schütt et al. (2017), PhysNet Unke and Meuwly (2019), SphereNet Liu et al. (2022) and GemNet Gasteiger et al. (2021), use internal features that are invariant under rotations. In models such as NewtonNet Haghighatlari et al. (2022), EGNN Satorras et al. (2021), and PaiNN Schütt et al. (2021), vector features built using relative atomic positions are also used on top of invariant features. In later equivariant MPNNs, such as Cormorant Anderson et al. (2019), NequIP Batzner et al. (2022), MACE Batatia et al. (2022a) exploit internal features based on irreducible representations of the E(3) symmetry group Geiger and Smidt (2022); Goodman and Wallach (2000).
Under the hood, in both ACE and E(3) equivariant MPNNs, the key process is the following: Equivariant features or messages of body order $\nu$ are encoded in spherical harmonics with degrees $l$. The invariant features are subsequently created by contracting the equivariant ones via the Clebsch-Gordan coefficients. These invariants are used to predict energies or other invariant physical quantities. Although the set of spherical harmonics forms a natural basis for the irreducible representations of the SO(3) group and thus is useful in operations involving rotational symmetries, it has drawbacks: the transformation of atomic coordinates to a spherical coordinate system can be an overhead although using Legendre polynomials alleviates this problem Drautz (2020), and the Clebsch-Gordan contraction is non-intuitive and cumbersome to use. As the degree $l$ and the body order $\nu$ increases, the matrix of Clebsch-Gordan coupling coefficients becomes very large ($\mathcal{O}(l+1)^{2\nu}$) but extremely sparse, and the number of the equivariant features increases exponentially when explicitly evaluated Darby et al. (2023); Nigam et al. (2020).
Here, we propose an alternative method that performs the expansion of atomic density as well as symmetrization of orientation-dependent features in Cartesian coordinates directly, circumventing spherical harmonics. The Cartesian atomic cluster expansion (CACE) has the same mathematical foundation as the ACE framework, so it also adopts the nice properties of ACE – body order and completeness in describing atomic environments Dusson et al. (2022). The CACE invariant features are low-dimensional and can be computed independently and efficiently. By further incorporating a low-dimensional embedding for chemical elements and message passing between neighboring atoms, we develop the CACE potential. We benchmark CACE on diverse systems, with emphasis on stability and extrapolation.
II Methods
Atomic graph
As illustrated in Fig. 1, each atomic structure is represented as a graph, with atoms as nodes, and directed edges that point from pairs of sender atoms to receiver atoms that are within a cutoff radius of $r_{\mathrm{cut}}$. The state of an atom $i$ is denoted by the tuple $\sigma_{i}=(\textbf{r}_{i},\boldsymbol{\theta}_{i},\textbf{h}_{i})$: $\textbf{r}_{i}$ is its Cartesian position, $\boldsymbol{\theta}_{i}$ encodes its chemical element $\alpha$, and $\textbf{h}_{i}$ is the learnable features. Later we will use the superscript, e.g. $\sigma_{i}^{(t)}$, to denote the atomic state in the message passing layer $t$, but for now we omit the layer-dependence indices for simplicity.
Figure 1: Schematic of the CACE potential. a-h show each step of the operation, and i illustrates the rule for making rotationally invariant features.
For both the sender node and the receiver node of an edge, the elements $\alpha_{i}$ are first mapped to vectors of lengths equal to the total number of elements $N_{\mathrm{element}}$ via one-hot embedding, and then the one-hot vectors are multiplied with a learnable weight matrix of size $N_{\mathrm{embedding}}\times N_{\mathrm{element}}$ outputting a learnable embedding $\boldsymbol{\theta}_{i}$ for each atom $i$ of length $N_{\mathrm{embedding}}$. Typically, $N_{\mathrm{embedding}}$ is selected to be a small number between 1 and 4. This continuous element embedding eliminates the unfavorable scaling of using discrete chemical element types when representing atomic environments, and also allows alchemical learning.
Edge basis
The edge basis, shown in Fig. 1b, is used to describe the spatial arrangement of an atom $j$ around the atom $i$ ($\vec{\mathbf{r}}_{ji}\equiv\mathbf{r}_{j}-\mathbf{r}_{i}$, $r_{ji}=|\vec{\mathbf{r}}_{ji}|$ and $\hat{\mathbf{r}}_{ji}=\vec{\mathbf{r}}_{ji}/r_{ji}$), and is formed by the product of a radial basis $R$, an angular basis $L$, and an edge-type basis $T$:
$\chi_{cn\mathbf{l}}(\sigma_{i},\sigma_{j})=T_{c}(\sigma_{i},\sigma_{j})R_{n,cl}(r_{ji})L_{\mathbf{l}}(\hat{\mathbf{r}}_{ji}).$ (1)
Eqn. (1) is general and fundamental for many MLIPs, and a systematic discussion of the design choices for each term can be found in Ref. Batatia et al. (2022b).
In CACE, the type of edge $T$ depends on the states of the two nodes it connects. In the initial layer, the edge states are only determined by the chemical elements of the two atoms $i$ and $j$. The edge is encoded using the flattened tensor product of the two node embedding vectors, $T=\boldsymbol{\theta}_{i}\otimes\boldsymbol{\theta}_{j}$, resulting in an edge feature of length $c=N_{\mathrm{embedding}}^{2}$. In a later message passing layer $t$, the edge features can depend on the node hidden features $\mathbf{h}_{i}^{(t)}$. Such edge encoding formed by the tensor product can be interpreted as the attention mechanism Vaswani et al. (2017) between the two atoms, with the sender and the receiver node embedding vectors being the query and the key.
The orientation-dependent angular basis is
$L_{\mathbf{l}}(\hat{\mathbf{r}}_{ji})=(x_{j}-x_{i})^{l_{x}}(y_{j}-y_{i})^{l_{y}}(z_{j}-z_{i})^{l_{z}},$ (2)
where $\mathbf{l}\equiv(l_{x},l_{y},l_{z})$ are the angular momentum numbers satisfying $l_{x}+l_{y}+l_{z}=l$, with $l$ being the total angular momentum. Such angular basis is just spherical harmonic functions in Cartesian basis. Indeed, one can easily transform the function of a certain angular momentum number $l$ in Cartesian and spherical harmonic bases by a simple matrix multiplication Altmann (1957).
$R_{n,cl}(r_{ji})$ is a set of radial basis consisting of $n$ functions, coupled with the total angular momentum $l$ and the edge channel $c$. This means that, for each $l$ and $c$ combination, there is a different set of $n$ radial functions. In CACE, the raw radial embedding block uses a set of trainable Bessel functions with different frequencies Schütt et al. (2021) multiplied by a smooth cutoff function. The output is then mixed using a linear transformation to form $R_{n,cl}$ for each $l$ and $k$.
Atom-centered basis ($A$-basis)
As shown in Fig. 1c, the atom-centered representation is made by summing over all the edges of a node,
$A_{i,cn\mathbf{l}}=\sum_{j\in\mathcal{N}(i)}\chi_{cn\mathbf{l}}(\sigma_{i},\sigma_{j}).$ (3)
This corresponds to a projection of the edge basis on the atomic density, which is commonly referred to as the “density trick” and was introduced originally to construct SOAP and bispectrum descriptors Bartók et al. (2013).
Symmetrized basis ($B$-basis)
The orientation-dependent $A$ features are symmetrized to get the invariant $B$ features of different body orders $\nu$ as sketched in Fig. 1c. For $l=0$, $A$ is already rotationally invariant, so the $\nu=1$ features $B^{(1)}_{i,cn0}=A_{i,cn\textbf{0}}$.
For $\nu\geq 2$, the general rule for constructing the symmetrized basis is illustrated in Fig. 1i. The key idea is that pairs of angular features need to have a shared factor of $\mathbf{l}$, when summed over to form invariants. The terms that only differ from others by permutations of the $l$-combinations are not considered. Importantly, if a combination contains any elements satisfying $\mathbf{l}=\mathbf{0}$ or $\mathbf{l}=\mathbf{l^{\prime}}$, the resulting invariant can be constructed using the product of lower-order invariants, and can thus be eliminated for redundancy. The CACE framework thus makes polynomially independent invariants, which are much more compact than ACE Drautz (2019) or MTPs Shapeev (2016), where the basis functions span the space of all invariants including linearly dependent terms Dusson et al. (2022). Other radial or edge terms that do not depend on $\mathbf{l}$ or only depend on the total angular momentum $l$ can be multiplied on top without breaking the rotational symmetry.
More specifically, for $\nu=2$ Zhang et al. (2021a),
$B_{i,cnl_{1}}^{(2)}=\sum_{\mathbf{l}_{1}}\mathcal{C}(\mathbf{l}_{1})A_{i,cn\mathbf{l}_{1}}^{2},$ (4)
summed over all $\mathbf{l}_{1}$ with $l_{1x}+l_{1y}+l_{1z}=l_{1}$. $\mathcal{C}(\mathbf{l})=\dfrac{l!}{l_{x}!l_{y}!l_{z}!}$ is the combinatorial prefactor. Doing the expansion explicitly, one can use the multinomial formula to show that Zhang et al. (2021a)
$B_{i,cnl_{1}}^{(2)}=\sum_{j,k\in\mathcal{N}(i)}T_{c}(\sigma_{i},\sigma_{j})R_{n,cl_{1}}(r_{ji})T_{c}(\sigma_{i},\sigma_{k})R_{n,cl_{1}}(r_{ki})\\ \times cos^{l_{1}}(\theta_{ijk}).$ (5)
So this term includes three-body contributions of $\vec{\mathbf{r}}_{ji}$ and $\vec{\mathbf{r}}_{ki}$ together with their enclosed angle $\theta_{ijk}$. For $\nu=3$,
$B_{i,cnl_{1}l_{2}}^{(3)}=\sum_{\mathbf{l}_{1},\mathbf{l}_{2}}\mathcal{C}(\mathbf{l}_{1})\mathcal{C}(\mathbf{l}_{2})A_{i,cn\mathbf{l}_{1}}A_{i,cn(\mathbf{l}_{1}+\mathbf{l}_{2})}A_{i,cn\mathbf{l}_{2}},$ (6)
which includes four-body interaction terms in the form $R_{n,cl_{1}}(r_{ji})R_{n,c(l_{1}+l_{2})}(r_{ki})R_{n,cl_{2}}(r_{hi})cos^{l_{1}}(\theta_{ijk})cos^{l_{2}}(\theta_{ikh})$. For $\nu=4$,
$B_{i,cnl_{1}l_{2}l_{3}}^{(4)}=\sum_{\mathbf{l}_{1},\mathbf{l}_{2},\mathbf{l}_{3}}\mathcal{C}(\mathbf{l}_{1})\mathcal{C}(\mathbf{l}_{2})\mathcal{C}(\mathbf{l}_{3})\\ \times A_{i,cn\mathbf{l}_{1}}A_{i,cn(\mathbf{l}_{1}+\mathbf{l}_{2})}A_{i,cn(\mathbf{l}_{2}+\mathbf{l}_{3})}A_{i,cn\mathbf{l}_{3}},$ (7)
which includes five-body contributions in the form $R_{n,cl_{1}}(r_{ji})R_{n,c(l_{1}+l_{2})}(r_{ki})R_{n,c(l_{2}+l_{3})}(r_{hi})R_{n,cl_{3}}(r_{gi})\\ cos^{l_{1}}(\theta_{ijk})cos^{l_{2}}(\theta_{ikh})cos^{l_{3}}(\theta_{ihg})$.
Based on the expansion, one can see that these $B^{(\nu)}$ terms are the same as the $B^{(\nu)}$ terms in the original ACE paper Drautz (2019), although many linearly or polynomially-dependent features are eliminated and the computation in the Cartesian space is much more straightforward. The $B$ basis is a complete basis of permutation and rotation invariant functions of the atomic environment Drautz (2019). In practice, one truncates the series up to some maximum values of $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$ to make the computation tractable. The dimension of the $B$ basis is $c\times n\times N_{L}$, where each factor comes from the number of edge channels, radial basis, and angular basis, respectively. Fig. 2 shows the total number of angular features $N_{L}$ up to certain $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$. The number of $N_{l}$ is compact even for $\nu_{\mathrm{max}}=5$. For most practical applications, $l_{\mathrm{max}}$ between 2 and 4 and $\nu_{\mathrm{max}}$ between 2 and 4 are sufficient for good accuracy when fitting CACE potentials.
Notice that in the CACE formulation, each invariant $B$ feature can be evaluated independently, and there is no need to precompute or store the high-dimensional many-body product basis in ACE Drautz (2019) or moment tensors in MTPs Shapeev (2016) even in the sparse form. The reason why this is efficient can be understood as it effectively exploits the extreme sparsity of the Clebsch-Gordan coupling matrix when forming invariants.
Figure 2: The total number of angular features $N_{L}$, for maximum values of $l_{\mathrm{max}}$ and $\nu_{\mathrm{max}}$.
Message passing
The framework above reproduces the original ACE, albeit formulated entirely in the Cartesian space. In ACE, the atomic energy of each atom is determined by the neighboring atoms within its cutoff radius $r_{\mathrm{cut}}$. On top of that, one can incorporate message passing, which iteratively propagates information along the edges, thereby communicating the local information and effectively enlarging the perceptive field of each atom. A systematic discussion on the design choices of message passing can be found in Refs. Batatia et al. (2022b); Bochkarev et al. (2024).
CACE uses two types of message passing (MP) mechanisms. The first type, denoted by the green arrow in Fig. 1e, communicates the $A$ information between neighboring atoms: the orientation-dependent message from atom $j$ to $i$ can be expressed as:
$m_{1,j\rightarrow i,cn\mathbf{l}}^{(t)}=F_{cnl}(r_{ji})A_{j,cn\mathbf{l}}^{(t)}$ (8)
where $F_{cnl}(r_{ji})$ is a filter function that takes the scalar value of the distance $r_{ji}$, and it can depend on edge channel, radial channel and total angular momentum. In practice, we use an exponential decay function with trainable exponent, to capture the physics that farther atoms have less effect, although other choices are possible.
The second MP mechanism, denoted by the two blue arrows in Fig. 1e, uses a recursive edge embedding scheme analogous to recursively embedded atom neural networks (REANN) Zhang et al. (2021a) and multilayer ACE (ml-ACE) Bochkarev et al. (2022):
$m_{2,j\rightarrow i,cn\mathbf{l}}^{(t)}=H_{cl}(B_{j}^{(t)})\chi_{cn\mathbf{l}},$ (9)
where $H_{cl}$ is a trainable scalar function that can optionally depend on edge channel and total angular momentum. We simply use a linear neural network (NN) layer for $H$.
The aggregated message is
$M_{i,cn\mathbf{l}}^{(t)}=\sum_{j\in\mathcal{N}(i)}m_{1,j\rightarrow i,kn\mathbf{l}}^{(t)}+m_{2,j\rightarrow i,kn\mathbf{l}}^{(t)}$ (10)
One can then update the $A$ representation of each node,
$A_{i,cn\mathbf{l}}^{(t+1)}=G(A_{i,cn\mathbf{l}}^{(t)},M_{i,cn\mathbf{l}}^{(t)}),$ (11)
and for $G$ we use a simple linear function. After obtaining the updated $A^{(t+1)}$, we again take the symmetrized $B$-basis as described before. It is easy to verify that these symmetrized representations are only dependent on the scalar values of the atomic distances and angles, and thus rotationally invariant. Note that one can make various design choices for $F$, $H$ and $G$ functions: they can be dependent on $c$, $n$ and $l$, without breaking the invariance at the symmetrization stage.
In practice, only one or no MP layer is typically used in CACE. This is because the use of higher-body-order atomic features and messages reduces the need for many MP layers, as in the case of MACE Batatia et al. (2022a).
Readout and output
After $T$ MP layers, all the resulting $B$ features at each layer are concatenated. These invariant features are compact; the number of the features can be computed as $T\times c\times n\times N_{L}$. Then, a multilayer perceptron (MLP) maps these invariant features to the target of the atomic energy of each atom $i$,
$E_{i}=MLP(B_{i}^{(0)}\oplus B_{i}^{(1)}\oplus\ldots\oplus B_{i}^{(T)}).$ (12)
In practice, for the readout function we use the sum of a linear NN and a multilayer NN, as the former preserves the body-ordered contributions and the latter considers the higher-ordered terms from the truncated series.
Finally, the total potential energy of the system is the sum of all the atomic energies of each atom, and the forces can readily be computed by taking the derivatives of the total energy with respect to atomic coordinates.
III Applications
III.1 Bulk water
As an example application on complex fluids, we demonstrate CACE on a dataset of 1,593 liquid water configurations, each of 64 molecules Cheng et al. (2019). The dataset was computed at the revPBE0-D3 level of density functional theory (DFT). 90% of the data are used in the training. We used a cutoff $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions and $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and no message passing ($T=0$) or one message passing layer ($T=1$). The fitted CACE models have relatively few parameters: 24,572 trainable parameters for the CACE water model with $T=0$, and 69,320 trainable parameters for the model with $T=1$. The root mean square errors (RMSE) of the energy per water molecule and forces on the validation set are presented in Table 1.
[…]
BPNN Cheng et al. (2019) DeepMD Wang et al. (2018) EANN Zhang et al. (2021b) linear ACE Witt et al. (2023) REANN Zhang et al. (2021a) NequIP Batzner et al. (2022) MACE Batatia et al. (2022a) CACE CACE
$6.35$ $6$ $6.2$ $5.5$ $6.2$, $T=3$ $4.5$, $T=3$ $6$ $5.5$, $T=0$ $5.5$, $T=1$
E 7.0 6.3 6.3 5.196 2.4 2.8 1.9 3.49 1.77
F 120 92 129 99 53.2 45 36.2 79 47
[…]
Refer to captionaRefer to caption Refer to captionb
Figure 3: Simulation results of water using CACE $T=1$ model. a Oxygen-oxygen radial distribution functions (RDF) at different temperatures and 1 g/mL computed via classical MD simulations in the NVT ensemble. The experimental O-O RDF at ambient conditions was obtained from Ref Skinner et al. (2014). b Mean squared displacement (MSD) from the liquid water simulations. Diffusivities ($D$) are shown in the legends.
For running molecular dynamics (MD) simulations using MLIPs, besides accuracy, a crucial requirement is the stability - the possibility to run long-enough trajectories with the system remaining in physically reasonable states and the simulation not exploding, even at conditions with few or no training data. To assess the simulation stability, we performed independent simulations of 300 ps in length for water at 1 g/mL at 300 K, 500 K, 1000 K, 1500 K, and 2000 K using the CACE $T=1$ model. The simulation cell contains 512 water molecules. The time step was 1 fs, and Nosé–Hoover thermostat was used. The upper panel of Fig. 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF). At 300 K, the computed O-O RDF is in excellent agreement with experiment from X-ray diffraction measurements Skinner et al. (2014). Note that the nuclear quantum effects have a slight destructuring effect in the O-O RDF, but the effect is quite small Cheng et al. (2019). The mean squared displacements (MSD) from these simulations are shown on the lower panel of Fig. 3. The diffusivity $D$, obtain by fitting a linear function to the MSD, at 300 K agrees well with the DFT MD result ($D_{\mathrm{DFT}}=2.67\pm 0.10\mathrm{\AA}^{2}/$ps) from Ref. Marsalek and Markland (2017). At high temperatures up to 2000 K, the MD simulations remained stable, which demonstrate the superior stability of the CACE potential. Note that we do not expect the result at the very high temperatures to be physically predictive, and this example is just to illustrate the stability of CACE. Such stability is not only necessary for converging statistical properties, but also allows the refinement of the potential: to iteratively improve the accuracy of the potential under certain conditions, one can perform active learning which collects snapshots from corresponding MD trajectories and use these snapshots for re-training.
III.2 Small molecules: ethanol and 3BPA
To demonstrate CACE on small organic molecules, we fitted to the ethanol sampled from DFT molecular dynamics simulations at 500 K, which is from the MD17 database Chmiela et al. (2017) that is often used for benchmark purposes. We used a cutoff $r_{\mathrm{cut}}=4$, 6 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=4$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=4$, and one message passing layer ($T=1$). The model was trained on 1,000 structures randomly selected from the dataset (90% train/10% validation split), and tested on another 10,000 random structures.
Refer to captionaRefer to caption Refer to captionb
Figure 4: The dihedral scan benchmark of the 3BPA molecule. a Molecular structure and three dihedral angles. b The dihedral potential energy landscape for $\beta=120^{\circ}$, as predicted by DFT and CACE. The white dots on the DFT potential energy surface correspond to all the training configurations that have $\beta$ between ${}^{\circ}$ and ${}^{\circ}$.
SchNet Schütt et al. (2017) DimeNet Gasteiger et al. (2019) sGDML Chmiela et al. (2019) PaiNN Schütt et al. (2021) SphereNet Liu et al. (2022) NewtonNet Haghighatlari et al. (2022) GemNet-T Gasteiger et al. (2021) NequIP Batzner et al. (2022) CACE
E 3.5 2.8 3.0 2.7 2.6 2.2 2.37
F 16.9 10.0 14.3 9.7 9.0 9.1 3.7 3.1 7.0
S $\uparrow$ 247 26 86 33 169 300 300
Table 2: Mean absolute error (MAE) of energy (E) and force (F) for ethanol in the MD-17 dataset, in units of meV and meV/Å, respectively, trained on 1000 reference configurations. The stability (S) in the unit of ps shows how long the simulation using the MLIP is physically realistic in MD of a maximum length of 300 ps at 500 K. Besides CACE, the errors are obtained from respective references, and the stability is from Ref. Fu et al. (2023).
Table 2 shows a comparison of the different MLIPs trained on energies and forces for the same MD17-ethanol dataset. The energy accuracy of CACE is similar to the NequIP model with $T=5$, although the CACE force error is larger. The stability (S) in Table 2 measures how long MD simulations at 500 K using the MLIP is physically sensical (no pathological behavior or enter physically prohibitive states) in runs with a maximum length of 300 ps, so 300 ps is the maximum score in this case Fu et al. (2023). The stability values for other MLIPs are from Ref. Fu et al. (2023) which used 10,000 training structures and a time step of 0.5 fs in MD. For CACE, we used 1,000 training structures and a time step of 1 fs, which is more stringent, but the MD had nevertheless remained stable. As observed in Refs. Fu et al. (2023); Stocker et al. (2022), force accuracy does not imply stability, and the latter is a key metric for MLIPs.
ACE Drautz (2019) sGDML Chmiela et al. (2019) GAP Bartók et al. (2010) ANI-pretrained Gao et al. (2020) ANI-2x Gao et al. (2020) NequIP Musaelian et al. (2023) MACE Batatia et al. (2022a) CACE
300 K, E 7.1 9.1 22.8 23.5 38.6 3.3 3.0 6.3
300 K, F 27.1 46.2 87.3 42.8 84.4 10.8 8.8 21.4
600 K, E 24.0 484.8 61.4 37.8 54.5 11.2 9.7 18.0
600 K, F 64.3 439.2 151.9 71.7 102.8 26.4 21.8 45.2
1200 K, E 85.3 774.5 166.8 76.8 88.8 38.5 29.8 58.0
1200 K, F 187.0 711.1 305.5 129.6 139.6 76.2 62.0 113.8
Table 3: Root mean square errors (RMSE) for energy (E) and force (F) on the 3BPA temperature transferability dataset, reported in units of meV and meV/Å. All models were trained on $T=300$ K data. Results of previous models except for NequIP and MACE are from Kovács et al. (2021). ANI-2x was trained on 8.9 million structures, and ANI-pretrained was pretrained and fine-tuned on this dataset.
As another example, we fitted to the 3BPA (3-(benzyloxy)pyridin-2-amine) dataset Kovács et al. (2021) consisting of 500 training structures at 300 K. Test data are from simulations at 300 K, 600 K, and 1200 K. We used $r_{\mathrm{cut}}=6$, 8 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=4$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=4$, and $T=1$. Table 3 shows the errors of different potentials when extrapolating to higher temperature. NequIP and MACE perform the best, while CACE is somewhere between them and ACE. We then compared the dihedral potential energy surface of the molecule. Fig. 4 shows the complex energy landscapes of the $\alpha$-$\gamma$ plane predicted by DFT and by CACE, for the case of $\beta=120^{\circ}$, the plane with the fewest training points. Nevertheless, the CACE landscape closely resembles the ground truth and is regular, more so than the other potentials benchmarked in Ref. Kovács et al. (2021) despite those were trained on a more diverse dataset including high temperature configurations.
III.3 25-element high-entropy alloys
High-entropy alloys, which are composed of several metallic elements, have unique mechanical, thermodynamics, and catalytic properties George et al. (2019). They are challenging to model due to the high number of elements involved, for which many MLIPs scale poorly with. Ref. Lopanitsyna et al. (2023) introduced 25 d-block transition metal HEA dataset that contains 25,630 distorted crystalline structures containing 36 or 48 on bcc or fcc lattices. Ref. Lopanitsyna et al. (2023) further performed an alchemical learning (AL) that explicit compresses chemical space, leading to a model that is both interpretable and very stable.
Figure 5: Benchmark on the HEA25 dataset. a Learning curves for different models. Both alchemical learning (AL) models are from Ref. Lopanitsyna et al. (2023). b The first two principal components (PCs) of the CACE embedding matrix. The periods are highlighted with orange, blue, and green lines. Interpolated positions for Re and Os are indicated with empty circles. c Energy comparison between DFT and CACE for a high temperature (5000 K) and a low temperature (300 K) trajectory. The inset shows example solid and melt configurations from the 300 K and the 5000 K trajectories, respectively. d Comparison between the substitution energy $E_{\mathrm{sub}}$, the potential energy difference between the original and the structure substituted with Re and Os element per substitution atom, computed using DFT and CACE. The inset shows the parity plot for the force components computed for those structures.
We took a test set containing 1,000 configurations, and used up to the rest of the 24,630 data points for training and validation. We used $r_{\mathrm{cut}}=4.5$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $T=1$, and $N_{\mathrm{embedding}}=4$ for aggressive alchemical compression.
In Fig. 5a, the CACE learning curve is compared to the ones from Ref. Lopanitsyna et al. (2023), including a 3B model that has an atomic energy baseline and three-body terms, and another 3B+NN model that further incorporates a full set of pair potentials and a non-linear term built on top of contracted power spectrum features. Overall, the CACE model achieved lower error and significantly more efficient learning for fewer data, and it has only 91,232 trainable parameters, compared to more than 160,000 in the 3B+NN model Lopanitsyna et al. (2023).
AL 3B+NN Lopanitsyna et al. (2023) PET Pozdnyakov and Ceriotti (2023) CACE
test, E 10 1.87 5.28
test, F 190 60.1 111
substitute, E 24 N.A. 8.5
substitute, F 373 N.A. 124
5000 K, E 48 152 15.3
5000 K, F 290 233
Table 4: Mean absolute error (MAE) of energy (E) and force (F) for the HEA25 dataset, reported in units of meV and meV/Å. There are three sets of benchmarks on test error, alchemical extrapolation to unseen Re and Os substitute atoms, and temperature transferability to 5000 K.
The errors of the models are presented in Table 4. A very recent architecture, Point Edge Transformer (PET) that utilizes local coordinate systems, achieved state-of-the-art performance in several benchmark systems and also has impressively low test error for the HEA25 dataset Pozdnyakov and Ceriotti (2023). However, as we discuss below, the PET potential has poor transferability.
Alchemical learning capability not only makes learning more efficient, but is also essential towards achieving a foundation model that is valid across the periodic table. CACE is able to perform alchemical learning. The learnable embedding $\boldsymbol{\theta}$ for each element type encodes its chemical information, and can be visualized to gain insights into the data-driven similarities. As the $N_{\mathrm{embedding}}=4$ in this case, we performed a principal component analysis (PCA) and plotted the first two principal component axes in Fig. 5b. The elements are arranged in a way that is strongly reminiscent of their placement in the d-block. This shows that the element embedding scheme is able to learn the nature of the periodic table in a data-driven way.
A stringent way to demonstrate alchemical learning is the extrapolation to unseen elements. In this case, Re and Os are absent from the training set. Ref. Lopanitsyna et al. (2023) provides a hold-out set containing 61 pairs of structures, each of 36 or 48 atoms. Each pair has one fcc or bcc structure with only the original 25 elements, and another structure with up to 11 atoms randomly substituted with Re and Os. We simply took the CACE model trained for the 25-element dataset, and obtained the chemical embedding of Re and Os by a linear interpolation between W and Ir, as illustrated by the empty circles in Fig. 5b. We also added atomic-energy baselines for Re and Os (obtained by fitting the residual energy to a two-parameter linear model that only depends on the Re and Os content). In Fig. 5d we show the parity plot for the substitution energy, defined as the potential energy difference between the original and the substituted structure per substitution atom. Impressively, the CACE error for those substituted structures with unseen elements are rather similar with the test error on the 25-element training set, and is lower than the AL method in Ref. Lopanitsyna et al. (2023). The force comparison, shown in the inset of Fig. 5d, shows the same trend. With the PET potential, it is not clear whether alchemical learning can be performed in a straightforward way.
The most impressive validation of the AL model in Ref. Lopanitsyna et al. (2023) is the extrapolation to configurations very far from the training set: At 5000 K, the systems generated from constant-pressure MD simulations with Monte Carlo element swaps have melted, but the AL potential that was trained exclusively on solid structures is still able to describe the energetics of the structures. This is a test that the PET model Pozdnyakov and Ceriotti (2023) did not perform well. To verify the generalizability of CACE across a wide temperature range, we recomputed the energies and forces of the structures from the 5000 K trajectory as well as another one at a low temperature of 300 K, and show the results in Fig. 5c and Table 4. The CACE errors are a bit smaller compared to Ref. Lopanitsyna et al. (2023), and quite modest compared to the training error.
IV Discussion and limitations
Atomic cluster expansion and equivariant MPNNs are probably both amongst the biggest breakthroughs in the field of MLIPs: the ACE formulation not only guarantees a complete basis for describing atomic environments, but also provides a framework for understanding and connecting different descriptors Drautz (2019); Dusson et al. (2022). E(3) equivariant MPNNs were able to achieve an unprecedented accuracy compared to the previous invariant architectures Batzner et al. (2022); Batatia et al. (2022a). It has been recognized that the spherical harmonics expansion and the subsequent Clebsch-Gordan tensor contraction is a bottleneck in the speed and the simplicity of ACE and these equivariant MPNNs. Clever methods have been developed to alleviate this, such as evaluate spherical harmonics in Cartesian space Drautz (2020); Lysogorskiy et al. (2021), absorbing Clebsch Gordan matrix in expansion coefficients during the evaluation stage Lysogorskiy et al. (2021), and using Gaunt coefficients instead of Clebsch-Gordan Luo et al. (2024). There have been other approaches that only operates in the Cartesian space, including NewtonNet Haghighatlari et al. (2022), EGNN Satorras et al. (2021), and PaiNN Schütt et al. (2021) that uses atomic displacement vectors in message passing, the PET based on local coordinate systems Pozdnyakov and Ceriotti (2023), and TensorNet that uses Cartesian tensors to represent pairwise displacements Simeon and De Fabritiis (2023). CACE is different from these approaches in that it can be directly mapped to ACE, so like ACE, it is also complete and body-ordered. The completeness helps the accuracy and the smoothness of the potential, particularly when there exists atomic environments that are degenerate using three-body descriptors or bispectrum Pozdnyakov et al. (2020); Nigam et al. (2023). The body order makes the potential more physically interpretable and may help extrapolation. The angular basis of CACE (Eqn. (2)) is the same as MTP Shapeev (2016), but CACE avoids computing moment tensors, and form a more compact symetric basis set via a simple combination rule that eliminates polynomially depedent features (Fig. 1i). As added advantages, the CACE descriptors are low-dimensional, efficient to evaluate, and are capable of alchemical learning.
A topic that tends to be overlooked in the development of MLIP methods is the stability and generalizability. In practice, a MLIP that is not stable in MD simulations is not useful. Nonetheless, stability and generalization are rarely considered in benchmarks, maybe due to that they are less straightforward to assess than energy and force errors. On the other hand, it has been shown that accuracy of the potential does not imply stability Fu et al. (2023); Stocker et al. (2022). The CACE potential shows high stability and extrapolatability in the datasets we benchmarked: stable MD simulations of water up to 2000 K, MD of ethanol, extrapolation to high temperature and dihedral scan of 3BPA, as well as generalization to the melt and unseen elements in high-entropy alloys. For water, the accuracy of CACE is on par with the most accurate potentials to date, NequIP Batzner et al. (2022) and MACE Batatia et al. (2022a). For small molecules, CACE is less accurate than NequIP or MACE but more accurate than the other recent MLIPs that we compared to. For all practical purposes, the accuracy of CACE for the small molecules is probably enough, considering the error is very small in absolute terms and is way below the chemical accuracy. Moreover, CACE achieved such accuracy with only one message passing layer and relatively few training parameters, and its accuracy can probably be improved with more message passing layers, higher $N_{\mathrm{embedding}}$, $\nu_{\mathrm{max}}$, $l_{\mathrm{max}}$, or a larger MLP at the readout.
We implemented the CACE potential using PyTorch, and the code is available in https://github.com/BingqingCheng/cace. With the current implementation, the training time of the water and the HEA25 datasets took about two days on a vintage GeForce GTX 1080 Ti GPU card, and the small molecules can be trained on a laptop within one or two days. The training should be much faster on state-of-the-art GPUs. For MD simulations of liquid water using the $T=1$ CACE water model on a single Nvidia A100 GPU, systems with 1,536 atoms, 5,184 atoms, 12,288 atoms, 24,000 atoms can run 24 ps per hour, 8 ps per hour, 3.5 ps per hour, 1.7 ps per hour respectively. The ability to run such large system sizes on a single GPU speaks to the low memory consumption of CACE. The implementation of the code can be further optimized: The computation of the node features and the symmetrization step may be compiled and be made faster. The training and the prediction are available on one GPU and can be made more parallelizable. The MD simulations can be performed in the ASE package, but the LAMMPS interface implementation is currently absent. As CACE has a moderate receptive field, so it is in principle possible to make it scale well in MD simulations.
Many engineering aspects of the CACE potential can be further fine-tuned. For example, dataset normalization and internal normalization can have a large effect on training efficiency and outcome Batatia et al. (2022b). Radial basis influences the learning efficiency Bigi et al. (2022), and its best choice is still an open question. Other choices of $F$, $H$ and $G$ functions in the message passing may be used. The training protocol can also influence the outcome.
Other possible future improvements of CACE include pretraining on large datasets before fine-tuning the model. It should be possible to make a foundation model that is generally valid across the periodic table Batatia et al. (2023). The alchemical learning capability of CACE may facilitate this.
V Conclusions
In summary, we propose a new scheme to preserve the rotational symmetry of atomic environments with different body orders, and the operation is entirely in the Cartesian coordinates. The transformation to the spherical harmonic basis, tensor product of atomic basis, and the usage of high-dimensional Clebsch-Gordan coupling matrix are akk avoided. The resulting symmeric features are low-dimensional and polynormially indepedent. As the symmetrization in the Cartesian coordinates is equivalent to the Clebsch-Gordan contraction, the scheme can be applied in other MLIPs that use the ACE framework (e.g. linear ACE Witt et al. (2023), SNAP, SOAP-GAP Bartók et al. (2010), Jacobi-Legendre cluster expansion Domina et al. (2023), ml-ACE Bochkarev et al. (2022), graph ACE Bochkarev et al. (2024)) , or E(3) equivariant message passing neural networks (e.g. NequIP Batzner et al. (2022), MACE Batatia et al. (2022a), BOTNet Batatia et al. (2022b), SEGNN Brandstetter et al. (2021), Allegro Musaelian et al. (2023)).
Data availability The water dataset is from from https://github.com/BingqingCheng/ab-initio-thermodynamics-of-water. MD17-ethanol is from http://www.sgdml.org/#datasets. BPA is from Ref. Kovács et al. (2021), downloaded from https://github.com/davkovacs/BOTNet-datasets. The HEA25 dataset is from Ref. Lopanitsyna et al. (2023), downloaded from  https://archive.materialscloud.org/record/2023.57.
The training scripts, trained CACE potentials, and MD input files are available at https://github.com/BingqingCheng/cacefit
Code availability The CACE package is publicly available at https://github.com/BingqingCheng/cace.